All questions please
101. The mass of the Earth is almost 6.0x104 kg, and the average radius is about 6400 km. A satellite with a mass of 70 Kg is orbiting at an altitude of 600 km. Calculate the acceleration with which the satellite is falling on Earth. a. 5.7x10 m/s b. 5.7x10 m/s c. 5.7x10° ms d. 5.7x10 m/ e. 5.7x10m/s 102. The mass of the Earth is almost 6.0x 10* kg, and the average radius is about 6400 km. A satellite with a mass of 70 Kg is orbiting at an altitude of 600 km. Calculate the acceleration with which the satellite is attracting the Earth. a. 6.7x10 m/s b. 6.7x10-1m/s c. 6.7x1016 m/s d. 6.7x1017 m/s2 e. 6.7x10-1 m/s 103. The first law of Kepler a. Defines the trajectory of the planed in a binary sun- planet system b. Specifies that all planets are moving around their sun on elliptical orbits. c. Indicates that the sun should be in one of the foci of the ellipses. d. Indicates that the other focus of the ellipses is empty e. All of the above. 104·The second law of Kepler says that The radius vector drawn from the Sun to a planet sweeps out equal areas in equal time intervals. Please check the correct statement. a. This is a new statement not included in the Newtons law of universal gravitation b. This is a generalization of the fact that the velocity of the planet engaged on a circular orbit is constant c. The generalization of the circular orbit to elliptical orbits implied that the speed of the plane is not constant and resulted in the second law of Kepler. d. Statements b and c are correct e. All statements (a, b, and c are correct) 105. Third Keppler law expresses the equivalence between the gravitational force and the centripetal force, and is defined as: a. The square of the orbital period of any planet is proportional to the cube of the axis of the elliptical orbit b. The orbital period of any planet is proportional to the cube of the semimajor axis of the elliptical orbit c. The square of the orbital period of any planet is proportional to the cube of the semimajor axis of the elliptical orbit d. The square of the orbital period of any planet is proportional to the cube of the semi-minor axis of the elliptical orbit e. The cube of the orbital period of any planet is proportional to the square of the semimajor axis of the elliptical orbit 106.The aphelion is: a. The point furthest away from the sun b. The point closest to the sun c. The average distance between the planet and the sun d. is a fixed point in the sky.
Answer #1

The given mass of Earth is M = 6.0 \small \times 1024 kg, the radius of Earth is R = 6400 km, mass of the satellite is m = 70 kg and the altitude is h = 600 km. Thus the acceleration with which satellite is falling on Earth can be calculated as,

  \small a=g\left ( 1-\frac{2h}{R} \right )

\small \Rightarrow a=9.8\left ( 1-\frac{2(600)}{6400} \right )

  \small \Rightarrow a=7.9625\: m/s^{2}